How many bit strings of length 11 have more 0s than 1s?


1
Expert's answer
2021-05-07T09:49:15-0400

The order of bits is not important since we are interested in the number of ones, not the order of ones. Thus, we need to use combinations.

"\displaystyle C(n,r) = \frac{n!}{r! (n-r)!}"

"n=11". There are more zeros than ones when less than 6 bits have ones.

"\displaystyle C(11,5) = \frac{11!}{6!\,5!} =462"

"\displaystyle C(11,4) = \frac{11!}{7!\,4!} =330"

"\displaystyle C(11,3) = \frac{11!}{9!\,3!} =165"

"\displaystyle C(11,2) = \frac{11!}{9!\,2!} =55"

"\displaystyle C(11,1) = \frac{11!}{10!\,1!} =11"

"C(11,0)=1"

So there are "462+330+165+55+11+1=1024" bit strings of length 11 that have more 0s than 1s.


Answer: 1024